milin ,hemant, and rashik have distributed profit of the firm of 1,89,000 in equal proportion . instead of distributing it in 1:2:4 proportion . to rectify this mistake what treatment will be given to the capital account
Answer by mathmate(429)

Question:
milin ,hemant, and rashik have distributed profit of the firm of 1,89,000 in equal proportion . instead of distributing it in 1:2:4 proportion . to rectify this mistake what treatment will be given to the capital account

Solution:
distributed share:
M:H:R = 1 : 1 : 1 = 7/21 : 7/21 : 7/21
Proper share among proprietors:
M:H:R = 1:2:4 = 1/7 : 2/7 : 4/7 = = 3/21 : 6/21 : 12/21
Corrections required: Proper share - distributed share
= 3/21-7/21 : 6/21-7/21 : 12/21-7/21
= -4/21 : -1/21 : 5/21 multiply by 1,89,000 [assumed correct, applicable to some countries]
= -36,000 : -9,000 : +45,000
So Milin has to reimburse 36,000, Hemant, 9000 while Rashik should receive 45,000.